OpenStudy (abbie):

what is an equation for the line that passes through points (-1, -4) and (1, 4)? i need help with this?!

OpenStudy (solomonzelman):

plug the points into slope formula

OpenStudy (abbie):

it says i have to write an equation in slope intercept form

OpenStudy (solomonzelman):

so find the slope of that equation first.

OpenStudy (abbie):

1/4

OpenStudy (solomonzelman):

Good! Now, I would graph it,

OpenStudy (solomonzelman):

so what is the y intercept?

OpenStudy (solomonzelman):

it [passes through the origin.

OpenStudy (abbie):

im not sure

OpenStudy (solomonzelman):

where on the line does x equal 0?

OpenStudy (abbie):

middle

OpenStudy (solomonzelman):

yes, it is the (0,0) so that would mean what?

OpenStudy (abbie):

so its one?

OpenStudy (solomonzelman):

the intercept? No!

OpenStudy (abbie):

oh sorry

OpenStudy (solomonzelman):

Look at the line, find the point on it which is on the 0 x-axis.

OpenStudy (anonymous):

Equation of line passing through two points is given as; \[y-y_1= \frac{y_1-y_2}{x_1-x_2} (x-x_1)\] Here (-1, -4) and (1, 4)? x1= -1, x2= 1 y1= -4, y2= 4 plugging these values in the above equation we find; \[y-(-4)= \frac{-4-4}{-1-1} [x-(-1)]\] \[\rightarrow y+4= \frac{-8}{-2} [x+1]\] \[\rightarrow y+4=4 (x+1) \rightarrow y+4=4x+4\] \[\rightarrow y=4x+4-4 \rightarrow y=4x+0 \rightarrow y=4x\] Hence y=4x is the required equation of the line. @abbie

OpenStudy (solomonzelman):

@dpasingh, we are guiding the asker, not giving out the work and answer!

OpenStudy (solomonzelman):

@abbie do you understand?

OpenStudy (abbie):

im getting confused

OpenStudy (solomonzelman):

Ok, where did you lose me?

OpenStudy (abbie):

im just going to skip this question

OpenStudy (solomonzelman):

no, you can do it.

OpenStudy (solomonzelman):

@abbie please don't skip it. Just tell me, I can redo it again.

OpenStudy (solomonzelman):

What don't you understand?

OpenStudy (abbie):

no thnx never mind i get it

OpenStudy (solomonzelman):

Nice to hear that! it is y=4x (as dpasingh said, so it is direct variation, in other words no y intercept.) can you go to your next question please?)